If an object is traveling at 37 m/s, how far will it travel in 10500 sec? Answer in km.
1
Expert's answer
2013-01-15T08:48:04-0500
To find distance, we have to multiply velocity by time of travel
S = v*t = 37m/s * 10500s = 388500 m = 388.5 km